OpenStudy (anonymous):

The function e(t)=4^8 * 4^t approximates the number of nematodes in a certain sample of fresh compost after t days. Find the initial number of nematodes when t = 0. How many nematodes are there after 3.5 days?

OpenStudy (johnweldon1993):

Well all we need to do is plug in 0 for 't'.....and then plug in 3.5 for 't' \[\large e(0) = 4^8 \times 4^0 = 4^8 = ?\]